# Federico Echenique and Farzad Pourbabaee

### 14 Jan 2024 v1

@@ 1 Introduction
First we consider individual welfare in a Walrasian equilibrium allocation. We quantify welfare by the probability that a random perturbation of an agent’s equilibrium consumption would provide a strict utility improvement: specifically, an improvement of “at least $\varepsilon$” over equilibrium consumption; meaning that the perturbation would remain an improvement if we were to shave off a fraction $\varepsilon$ from consumption. Our first result is actually more general, and quantifies the perturbations that could make at least one agent better off by at least $\varepsilon$, but we focus on a simple special case in this introduction. Now, it is obvious that an agent’s consumption can be improved whenever their preferences are monotonic, and the probability of an improvement by perturbation is always strictly positive. Our first result says that the probability that a perturbation is an individual $\varepsilon$-improvement converges (exponentially) to zero as the number of states grows.
To understand our first result, consider the situation of an individual agent $i$ in equilibrium. The usual consumption-maximization diagram is reproduced in Figure 1 on the left. In the figure, there are two states of the world ($d=2$), and an agent’s budget set is depicted in gray. The budget is defined by an equilibrium price $p$, and the agent’s endowment $\omega_{i}$. The picture shows the agent’s equilibrium consumption $f_{i}$. The agent’s indifference curve, tangent to the budget line at $f_{i}$, is depicted as a dashed curve. The upper contour set is the region to the northeast of the indifference curve that contains all consumptions that are preferred to $f_{i}$.
@@
We wish to assess the size of the upper contour set at $f_{i}$: what is the scope for improving $i$’s welfare at the equilibrium allocation $f$? By virtue of the monotonicity of $i$’s preference, the upper contour set is, of course, infinite. It contains all the consumption vectors that are larger, state by state, than $f$. So we consider its size relative to a ball that is centered at $f_{i}$. Specifically, we measure the size of the dotted region on the right panel of Figure 1. Think of the ball as a set of perturbations of $f_{i}$. [^1] The ball drawn in the figure includes negative consumption, but its radius may be taken to be arbitrarily small so as to avoid negative consumption, if so desired. We may also restrict attention to consumption that is unaffordable at equilibrium prices (the region to the northeast of the budget line) without changing the essential message of our result. We wish to quantify how often such perturbations would lead to a utility improvement for $i$. Key to our result is that the improvement should be at least by $\varepsilon$; for simplicity the figure only depicts the strictly better bundles (the case of $\varepsilon=0$).
@@
The consumer would be better off with more consumption in every state, but they would also be willing to accept a tradeoff of more consumption in one state than in another. Of course, the tradeoff would have to be at terms of trade that are more favorable than equilibrium prices. But the figure shows that there is a significant probability that a perturbation would leave the agent strictly better off, and the upper contour contains much more than the bundles that are larger state-by-state. So does the dotted region. Our first theorem says that, when the number of states is $d$, the probability of the dotted region, calculated for $\varepsilon$ improvements, is bounded above by $\mathrm{e}^{-\varepsilon^{2}d/8}$. In particular, it converges rapidly to zero as $d$ grows large.
@@
To put this bound in perspective, note that the fraction of bundles that are larger, state-by-state, than $f_{i}$ also converges exponentially to zero. So our result says that, despite the potential for utility improvements by trading off more consumption in some states for less consumption in other, the scope for such potential improvement shrinks when the number of states is large. In a sense, the curvature of utility does not matter in high-dimensional trade. Our bound of $\mathrm{e}^{-\varepsilon^{2}d/8}$ holds for any monotonic and quasi-concave utility function.
@@
Our second main result concerns collective welfare in a Pareto optimal allocation. The exercise is similar to our first result, but relies on redistributing aggregate consumption so as to make all agents better off. For our second result, we consider an economy without aggregate uncertainty, and fix an allocation that is Pareto optimal. This time we consider a perturbation of aggregate consumption that can be redistributed to make all agents better off by at least $\varepsilon$ — meaning again a utility improvement that remains, even after shaving off a fraction $\varepsilon$ of consumption. We show that a perturbation to aggregate consumption that can be redistributed to improve all agents by at least $\varepsilon$ has vanishingly small probability.
@@
Pareto optimality refers to efficiency with a fixed aggregate endowment, or aggregate consumption. One cannot make all agents better off without changing the aggregate consumption. Our result says, when the state space is large, the scope for collective welfare improvements by a given margin $\varepsilon$ is limited, even by changing aggregate consumption. Again it is useful to compare our bound with the fraction of aggregate allocations that are larger than the endowment, state-by-state. The fraction larger allocations shrinks exponentially to zero, as does (according to our results), the allocation that afford a collective improvement.
@@
Our second result can be interpreted in light of Debreu’s coefficient of resource utilization (Debreu, 1951). If we consider an allocation that is not Pareto optimal, then Debreu’s coefficient of resource utilization measures the degree of inefficiency inherent in the allocation. We can again study the probability of an improving perturbation to aggregate consumption that makes agents better off (not by $\varepsilon$ this time, simply strictly better off). Our second main result implies that this probability shrinks to zero exponentially as the number of states grows large. If Debreu’s coefficient of resource utilization is $\mathrm{CRU}$, and Pareto inefficiency implies that $\mathrm{CRU}<1$, then the probability is bounded above by $\mathrm{e}^{-(1-\mathrm{CRU})^{2}d/8}$. [^2] Debreu (1951) proposes $\mathrm{CRU}\in(0,1]$ as a measure of inefficiency of an allocation, and shows that $\mathrm{CRU}=1$ if and only if the allocation is Pareto optimal.
@@
Our third main result is not explicitly about welfare, but instead about the attitudes towards uncertainty by agents who engage in efficient risk sharing. We consider the same setting of economic exchange with no aggregate uncertainty, as in our second result; but we strengthen the assumptions on preferences to focus on utilities with multiple priors. Think, to fix ideas, on the max-min expected utility preferences of Gilboa and Schmeidler (1989). Now we prove that, if the risk sharing agreement between the two agents can be improved in the Pareto sense, then at least one of the two must have a small set of prior beliefs. The stronger is the level of Pareto improvement, the smaller will be the size of the prior beliefs set. This means that at least one of the two agents must, in some sense, be close to being ambiguity neutral.
@@
Our results rely on the study of concentration of measures, and specifically on the implication of an Isoperimetric inequality to the separation of convex sets. Key is the role of $\varepsilon$, or of $\mathrm{CRU}$, in our discussion above. These magnitudes provide a certain “padding,” or quantitative bound, on the degree of separation of the relevant convex sets. The phenomenon of concentration of measure then has strong consequences for the volume of one of the sets being separated, and in our economic models we can determine which of the sets must have small volume (except for our third application, which is more subtle). Given its role in our results, we discuss the magnitude of $\varepsilon$ in Section 4.2.
@@ 2.1 Notations and Conventions
Let $A$ be a subset of a finite-dimensional normed vector space $(\mathbb{R}^{m},\lVert\cdot\rVert)$. The distance of a point $x\in\mathbb{R}^{m}$ from $A$ is defined by $\mathsf{dist}(x,A)\coloneqq\inf_{a\in A}\lVert x-a\rVert$. By default, the norms used in the paper are the Euclidean $\ell_{2}$ norm. When a particular $p$-norm is used, we refer to the distance function by $\mathsf{dist}_{p}$ and the norm by $\lVert\cdot\rVert_{p}$. For the $\ell_{2}$ norm, we omit the subscript 2.
@@
Given a measurable subset $A\subseteq\mathbb{R}^{m}$, we denote its Euclidean volume by $\mathsf{Vol}(A)$ that is equal to the Lebesgue integral of the indicator function of $A$ relative to the affine hull of $A$. For example, if $A$ is a $m-1$ dimensional surface in $\mathbb{R}^{m}$, then $\mathsf{Vol}(A)$ refers to the surface area of $A$, as opposed to its $m$ dimensional volume (which is zero).
@@ 2.2 Preferences and Uncertainty
We consider a setting with uncertainty. [^3] Using the reinterpretation of commodities in Debreu (1959), our results have implications for many other economic environments. For example, the textbook version of the model assumes that consumption is in units of physically distinct goods. Let $S$ be a finite set of possible states of the world, and $d$ be the number of states of the world, i.e., $d\coloneqq|S|$.
@@
  • (Continuity): The upper and lower contour sets are closed. That is for every $f\in\mathbb{R}_{+}^{d}$, the sets $\{g:g\succeq f\}$ and $\{g:f\succeq g\}$ are closed.
@@
Many well-known models in the theory of choice under uncertainty are special cases of convex preferences. Examples are risk-averse subjective expected utility (SEU), max-min expected utility (MEU: Gilboa and Schmeidler, 1989), Multiplier preferences (Hansen and Sargent, 2001), variational preferences (Maccheroni et al., 2006), and smooth ambiguity aversion (Klibanoff et al., 2005).
@@
By convexity and continuity of preferences the approximate upper contour set is a convex and open subset. Its convexity is of particular importance to us, as our arguments often rely on the hyperplane separation theorem.
@@ 2.3 Exchange Economies
An exchange economy is a mapping $\mathcal{E}:I\to\mathcal{P}\times\mathbb{R}_{+}^{d}$, where $\mathcal{E}(i)=(\succeq_{i},\omega_{i})$. Each agent $i\in I$ is described by a preference relation $\succeq_{i}$ on $\mathbb{R}_{+}^{d}$, as well as an endowment vector $\omega_{i}\in\mathbb{R}_{+}^{d}$. In an exchange economy, we use $\mathcal{U}_{i}^{(\varepsilon)}$ to denote the upper contour set $\mathcal{U}_{\succeq_{i}}^{(\varepsilon)}$.
@@
An allocation $f=\{f_{i}:i\in I\}$ is called a Walrasian equilibrium for the exchange economy $\mathcal{E}$, if there exists a price vector $p\in\mathbb{R}^{d}$ such that $g_{i}\succ_{i}f_{i}$ implies that $p\,\cdot\,g_{i}>p\,\cdot\,\omega_{i}$, and $p\,\cdot\,f_{i}=p\,\cdot\,\omega_{i}$, for every $i\in I$.
An exchange economy is convex if each preference relation $\succeq_{i}$ is convex, i.e., $\succeq_{i}\in\mathcal{C}$. Convexity is required for the basic theory of general equilibrium: existence of Walrasian equilibrium, as well as the second welfare theorem, relies on convex preferences. All of our results, except Proposition 1 and Theorem 3, will also rely on convexity
When $s\mapsto\sum_{i\in I}\omega_{i}(s)$ is constant, we say that $\mathcal{E}$ exhibits no aggregate uncertainty. The aggregate endowment is then same across all states of the world, i.e., $\omega=(\bar{\omega},\ldots,\bar{\omega})$.
@@ 3 Main Results
We proceed to state our main results, leaving a broader discussion our findings to Section 4, and an overview of the methodology behind our proofs to Section 5.
3.1 Optimality and Equilibrium
Our first result is a generalization of the property that we discussed in the introduction. It entails the following exercise: Fix a Walrasian equilibrium allocation, and draw a random perturbation from a ball of radius $r$. Consider the event that this random perturbation, when applied to any agent’s equilibrium consumption, results in contingent consumption that is better by at least $\varepsilon$ than their equilibrium consumption. The resulting event, that at least one agent is made better off, has vanishingly small probability.
@@ 3.1 Optimality and Equilibrium
Let $\mathcal{E}$ be a convex exchange economy. Suppose that there exists $\tau>0$ such that $\omega_{i}\geq\tau\bm{1}$ for all $i\in I$, and let $f$ be a Walrasian equilibrium allocation. Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. Then for every $\varepsilon>0$,
@@
The proof of this theorem is in Section 5, as is the proof of the rest of the results we provide in this section. To unpack the statement of the theorem, imagine an economy with two consumers, $I=\{1,2\}$. If a random perturbation $z\in\mathbb{B}(r)$ is also in $(\mathcal{U}_{1}^{(\varepsilon)}(f_{1})-f_{1}\})\cup(\mathcal{U}_{2}^{(\varepsilon)}(f_{2})-\{f_{2}\})$, then it means that either $(1-\varepsilon)(z+f_{1})$ provides a strict welfare improvement for agent 1 over $f_{1}$, or that $(1-\varepsilon)(z+f_{2})$ does this for agent 2 over $f_{2}$ (or that both things are true). If we take $\tau=r$ and consider a 10% welfare improvement ($\varepsilon=0.1$), then the probability of making at least one agent better off is at most $e^{-d/800}$. If $d$ is, say, the number of stocks trading on the NASDAQ Exchange, then this bound is about 1 %.
@@
By the monotonicity of preferences, $z>0$, a perturbation that leaves an agent with greater consumption state-by-state, is always a welfare improvement for any agent. Such a perturbation has probability $1/2^{d}$, which also decreases exponentially in the number of states. The “curvature” of agents’ preferences means, however, that many other perturbations should also result in a welfare improvement — recall the dotted area in Figure 1. So the conclusion in Theorem 1 may be surprising; it says that when the improvement is by a margin $\varepsilon$, the scope for welfare improvements shrinks exponentially in $d$ regardless of the shape of agents’ preferences, and even if we only ask for an improvement in one agent’s welfare.
As a consequence of Theorem 1, we obtain the statement that we discussed at length in the introduction:
@@
We should emphasize that we interpret volume as a uniform probability measure on the ball of radius $r$, so that the statement in Corollary 1 corresponds to the informal description we provided in the introduction. The volume of the ball itself shrinks exponentially in $d$, so it is important to condition on the ball.
@@
Our second result considers welfare improvements outside of equilibrium. An equilibrium allocation must be Pareto optimal, but the Pareto set is strictly larger. So we ask about the possibility of a collective improvement after a change in aggregate consumption, starting from an allocation that is a weak Pareto optimum.

### 14 Feb 2024 v2

@@ 1 Introduction
First we consider individual welfare in a Walrasian equilibrium allocation. We quantify welfare by the probability that a random perturbation of an agent’s equilibrium consumption would provide a strict utility improvement: specifically, an improvement of “at least $\varepsilon$” over equilibrium consumption; meaning that the perturbation would remain an improvement if we were to shave off a small fraction $\varepsilon>0$ from the agent’s consumption. (Our first result is actually more general, and quantifies the perturbations that could make at least one agent better off by at least $\varepsilon$, but we focus on a simple special case here in the introduction.) Now, it is obvious that an agent’s consumption can be improved whenever their preferences are monotonic, and the probability of an improvement by perturbation is always strictly positive. Our first result says that the probability that a perturbation is an individual $\varepsilon$-improvement converges (exponentially) to zero as the number of states grows.
To understand our first result, consider the situation of an individual agent $i$ in equilibrium. The usual consumption-maximization diagram is reproduced in Figure 1 on the left. In the figure, there are two states of the world ($d=2$), and an agent’s budget set is depicted in gray. The budget is defined by an equilibrium price $p$, and by the agent’s endowment $\omega_{i}$. The picture shows the agent’s equilibrium consumption $f_{i}$. The agent’s indifference curve, tangent to the budget line at $f_{i}$, is depicted as a dashed curve. The upper contour set is the region to the northeast of the indifference curve that contains all consumptions that are preferred to $f_{i}$.
@@
We wish to assess the size of the upper contour set at $f_{i}$: what is the scope for improving $i$’s welfare at the equilibrium allocation $f$? By virtue of the monotonicity of $i$’s preference, the upper contour set is, of course, infinite. It contains all the consumption vectors that are larger, state by state, than $f$, and more. So we consider the size of the upper contour set relative to a ball that is centered at $f_{i}$. Specifically, we measure the size of the dotted region on the right panel of Figure 1. Think of the ball as a set of perturbations of $f_{i}$. [^1] The ball drawn in the figure includes negative consumption, but its radius may be taken to be arbitrarily small so as to avoid negative consumption, if so desired. We may also restrict attention to consumption that is unaffordable at equilibrium prices (the region to the northeast of the budget line) without changing the essence of our results. Indeed, in the latter case, we could draw the figure so that the dotted region has probability close to one; still, with large $d$, its probability vanishes. We wish to quantify how often such perturbations would lead to a utility improvement for $i$. Key to our result is that the improvement should be at least by $\varepsilon$; for simplicity the figure only depicts the strictly better bundles (the case of $\varepsilon=0$).
@@
The consumer would be better off with more consumption in every state, but they would also be willing to accept a tradeoff of more consumption in one state than in another. Of course, the tradeoffs would have to be at terms of trade that are more favorable than equilibrium prices, and guided by the curvature of the agent’s utility function. But the figure shows that there is a significant probability that a perturbation would leave the agent strictly better off, and the upper contour contains much more than the bundles that are larger state-by-state. So does the dotted region. Our first theorem says that, when the number of states is $d$, the probability of the dotted region, calculated for $\varepsilon$ improvements, is bounded above by $\mathrm{e}^{-\varepsilon^{2}d/8}$. In particular, it converges rapidly to zero as $d$ grows large.
@@
To put this bound in perspective, note that the fraction of bundles that are larger, state-by-state, than $f_{i}$ also converges exponentially to zero (it is $(1/2^{d})$). So our result says that, despite the potential for utility improvements by trading off more consumption in some states for less consumption in other, the scope for such potential improvements shrinks when the number of states is large. In a sense, the curvature of utility does not matter in high-dimensional trade. Our bound of $\mathrm{e}^{-\varepsilon^{2}d/8}$ holds for any monotonic utility function.
@@
Our second main result concerns collective welfare in a Pareto optimal allocation. The exercise is similar to our first result, but relies on redistributing aggregate consumption so as to make all agents better off. For our second result, we consider an economy without aggregate uncertainty, and fix an allocation that is Pareto optimal. This time we consider a perturbation of aggregate consumption that can be redistributed to make all agents better off by at least $\varepsilon$ — meaning, again, a utility improvement that remains, even after shaving off a fraction $\varepsilon$ of consumption. We show that a perturbation to aggregate consumption that can be redistributed to improve all agents by at least $\varepsilon$ has vanishingly small probability.
@@
Pareto optimality refers to efficiency with a fixed aggregate endowment, or aggregate consumption. One cannot make all agents better off without changing aggregate consumption. Our result says, when the state space is large, the scope for collective welfare improvements by a given margin $\varepsilon$ is limited, even by changing aggregate consumption. Again it is useful to compare our bound with the fraction of aggregate allocations that are larger than the endowment, state-by-state. The fraction larger allocations shrinks exponentially to zero, as does (according to our results), the allocation that afford a collective improvement.
@@
Our second result can be interpreted in light of Debreu’s coefficient of resource utilization (Debreu, 1951). If we consider an allocation that is not Pareto optimal, then Debreu’s coefficient of resource utilization measures the degree of inefficiency inherent in the allocation. We may again study the probability of an improving perturbation to aggregate consumption that makes agents better off (not by $\varepsilon$ this time, simply strictly better off). Our second main result implies that this probability shrinks to zero exponentially as the number of states grows large. If Debreu’s coefficient of resource utilization is $\mathrm{CRU}$, and Pareto inefficiency implies that $\mathrm{CRU}<1$, then the probability is bounded above by $\mathrm{e}^{-(1-\mathrm{CRU})^{2}d/8}$. [^2] Debreu (1951) proposes $\mathrm{CRU}\in(0,1]$ as a measure of inefficiency of an allocation, and shows that $\mathrm{CRU}=1$ if and only if the allocation is Pareto optimal.
@@
Our third main result is not explicitly about welfare, but instead about the attitudes towards uncertainty by agents who engage in efficient risk sharing. We consider the same setting of economic exchange with no aggregate uncertainty, as in our second result; but we strengthen the assumptions on preferences to focus on utilities with multiple priors. Think, to fix ideas, on the max-min expected utility preferences of Gilboa and Schmeidler (1989) and of a setting with two agents. Now we prove that, if the risk sharing agreement between the two agents can be improved in the Pareto sense, then at least one of the two must have a small set of prior beliefs. The stronger is the level of Pareto improvement, the smaller will be the size of the prior beliefs set. This means that at least one of the two agents must, in some sense, be close to being ambiguity neutral.
@@
Our results rely on the study of concentration of measures, and specifically on the implication of an Isoperimetric inequality to the separation of convex sets. Key is the role of $\varepsilon$, or of $\mathrm{CRU}$, in our discussion above. [^3] Given its role in our results, we discuss the magnitude of $\varepsilon$ in Section 4.2. These magnitudes provide a certain “padding,” or quantitative bound, on the degree of separation of the relevant convex sets. The phenomenon of concentration of measure then has strong consequences for the volume of one of the sets being separated, and in our economic models we can determine which of the sets must have small volume (except for our third application, which is more subtle). Indeed the irrelevance of the curvature in agents’ indifference curves, or upper contour sets, is a familiar phenomenon in the literature on isoperimetric inequalities in high dimension.
@@ 2.1 Notations and Conventions
Let $A$ be a subset of a finite-dimensional normed vector space $(\mathbb{R}^{m},\lVert\cdot\rVert)$. The distance of a point $x\in\mathbb{R}^{m}$ from $A$ is defined by $\mathsf{dist}(x,A)\coloneqq\inf_{a\in A}\lVert x-a\rVert$. By default, the norms used in the paper are the Euclidean $\ell_{2}$ norm. When a particular $p$-norm is used, we refer to the distance function by $\mathsf{dist}_{p}$ and the norm by $\lVert\cdot\rVert_{p}$. For the $\ell_{2}$ norm, we ussualy omit the subscript 2.
@@
Given a measurable subset $A\subseteq\mathbb{R}^{m}$, we denote its Euclidean volume by $\mathsf{Vol}(A)$ that is equal to the Lebesgue integral of the indicator function of $A$ relative to the affine hull of $A$. For example, if $A$ is a $(m-1)$-dimensional surface in $\mathbb{R}^{m}$, then $\mathsf{Vol}(A)$ refers to the surface area of $A$, as opposed to its $m$-dimensional volume (which is zero).
@@ 2.2 Preferences and Uncertainty
We consider a setting with uncertainty. [^4] Using the reinterpretation of commodities in Debreu (1959), our results have implications for many other economic environments. For example, to the textbook version of the model which assumes that consumption is in units of physically distinct goods. Let $S$ be a finite set of possible states of the world, and $d$ be the number of states of the world; $d\coloneqq|S|$.
@@
  • (Continuity): The upper and lower contour sets are closed.
    That is for every $f\in\mathbb{R}_{+}^{d}$, the sets $\{g:g\succeq f\}$ and $\{g:f\succeq g\}$ are closed subsets of $\mathbb{R}^{d}$.
@@
Many well-known models in the theory of choice under uncertainty are special cases of convex preferences. Examples are risk-averse subjective expected utility (SEU), max-min expected utility (MEU: see Gilboa and Schmeidler, 1989), Multiplier preferences (Hansen and Sargent, 2001), variational preferences (Maccheroni et al., 2006), and smooth ambiguity aversion (Klibanoff et al., 2005).
@@
By convexity and continuity of preferences, the approximate upper contour set is a convex and open set. Its convexity is of particular importance to us, as our arguments often rely on the hyperplane separation theorem.
@@ 2.3 Exchange Economies
An exchange economy is a mapping $\mathcal{E}\colon I\to\mathcal{P}\times\mathbb{R}_{+}^{d}$, where $\mathcal{E}(i)=(\succeq_{i},\omega_{i})$. Each agent $i\in I$ is described by a preference relation $\succeq_{i}$ on $\mathbb{R}_{+}^{d}$, as well as an endowment vector $\omega_{i}\in\mathbb{R}_{+}^{d}$. In an exchange economy, we use $\mathcal{U}_{i}^{(\varepsilon)}$ to denote the upper contour set $\mathcal{U}_{\succeq_{i}}^{(\varepsilon)}$.
@@
An allocation $f=\{f_{i}:i\in I\}$ is called a Walrasian equilibrium allocation for the exchange economy $\mathcal{E}$, if there exists a price vector $p\in\mathbb{R}^{d}$ such that $g_{i}\succ_{i}f_{i}$ implies that $p\,\cdot\,g_{i}>p\,\cdot\,\omega_{i}$, and $p\,\cdot\,f_{i}=p\,\cdot\,\omega_{i}$, for every $i\in I$.
An exchange economy $\mathcal{E}$ is convex if each preference relation $\succeq_{i}$ is convex, i.e., $\mathcal{E}\colon I\to\mathcal{C}\times\mathsf{E}^{d}_{+}$. Convexity is required for the basic theory of general equilibrium: existence of Walrasian equilibrium, as well as the second welfare theorem, relies on convex preferences. Our Theorems 2 and 3, will also rely on convexity of preferences.
When $s\mapsto\sum_{i\in I}\omega_{i}(s)$ is constant, we say that $\mathcal{E}$ exhibits no aggregate uncertainty. The aggregate endowment is then same across all states of the world, i.e., $\omega=(\bar{\omega},\ldots,\bar{\omega})$. The assumption of no aggregate uncertainty is common in the study of risk sharing: see for example Allen and Gale (1988) or, more closely related to our paper, Billot et al. (2000). Our Theorem 2 requires that there be no aggregate uncertainty.
@@ 3 Main Results
We proceed to state our main results, leaving a broader discussion of our findings to Section 4. The proofs of our results, including an overview of the methodology behind the proofs, can be found in Section 5.
3.1 Walrasian Equilibrium
Our first result is a generalization of the property that we have discussed in the introduction. It entails the following exercise: Fix a Walrasian equilibrium allocation, and draw a random perturbation from a ball of radius $r$. Consider the event that this random perturbation, when applied to any agent’s equilibrium consumption, results in contingent consumption that is better by at least $\varepsilon$ than their equilibrium consumption. The resulting event, that at least one agent is made better off, has vanishingly small probability.
@@ 3.1 Walrasian Equilibrium
Let $\mathcal{E}\colon I\to\mathcal{P}\times\mathbb{R}_{+}^{d}$ be an exchange economy. Suppose that there exists $\tau>0$ such that $\omega_{i}\geq\tau\bm{1}$ for all $i\in I$, and let $f$ be a Walrasian equilibrium allocation. Fix $r>0$ and let $z\sim\mathsf{P}^{r}$. Then for every $\varepsilon>0$,
@@
The proof of this theorem is in Section 5, as is the proof of the rest of the results we provide in this section. To unpack the statement of the theorem, imagine an economy with two consumers, $I=\{1,2\}$. If a random perturbation $z\in\mathbb{B}(r)$ is also in $(\mathcal{U}_{1}^{(\varepsilon)}(f_{1})-f_{1}\})\cup(\mathcal{U}_{2}^{(\varepsilon)}(f_{2})-\{f_{2}\})$, then it means that either $(1-\varepsilon)(z+f_{1})$ provides a strict welfare improvement for agent 1 over $f_{1}$, or that $(1-\varepsilon)(z+f_{2})$ does this for agent 2 over $f_{2}$ (or that both things are true). If we take $\tau=r$ and consider a 10% welfare improvement ($\varepsilon=0.1$), then the probability of making at least one agent better off is at most $e^{-d/800}$. Financial economists often take the state space to equal the returns to a set of assets that are traded in the market: Hence, if we take $d$ to be at least the number of stocks, say, trading on the NASDAQ Exchange, then our bound from Theorem 1 is about 1 %.
@@
By the monotonicity of preferences, $z>0$, a perturbation that leaves an agent with greater consumption state-by-state, is always a welfare improvement for any agent. Such a perturbation has probability $1/2^{d}$, which also decreases exponentially in the number of states. The “curvature” of agents’ preferences means, however, that many other perturbations should also result in a welfare improvement — recall the dotted area in Figure 1. So the conclusion in Theorem 1 may be surprising; it says that when the improvement is by a margin $\varepsilon$, the scope for welfare improvements shrinks exponentially in $d$ regardless of the shape of agents’ preferences, and even if we only ask for improvement in one agent’s welfare.
Finally, we should emphasize that Theorem 1 bounds the probability of an improvement to at least one agent, while the bound is independent of the number of agents. We may then consider an economy with many agents and many states, and conclude that it is very unlikely that a given perturbation could make even one agent better of by at least $\varepsilon$. Now if we do focus on a single agent, as we did in the introduction, then we obtain the statement that was discussed at length in the introduction:
@@
We should emphasize that we interpret volume as a uniform probability measure on the ball of radius $r$, so that the statement in Corollary 1 corresponds to the informal description we provided in the introduction. The volume of the ball shrinks itself exponentially in $d$, so it is important to condition on the ball.
@@
3.2 Optimality and resource utilization
Our second result considers collective welfare improvements, and concerns outcomes that may not be Walrasian equilibria. An equilibrium allocation must be Pareto optimal, but the Pareto set is strictly larger. So we ask about the possibility of a collective improvement after a change in aggregate consumption, starting from an allocation that is a weak Pareto optimum.